OpenStudy (anonymous):

After the expression (x^10)^3/5 is simplified as much as possible, x is raised to what exponent?

OpenStudy (aum):

\[ \Large (x^m)^n = x^{m \times n} \\ \Large (x^{10})^{3/5} = x^{10 \times 3/5} = x^{30/5} = x^6 \\ \]